# Heisenberg limit in phase measurements: the threshold detection approach

### v1

@@ Abstract
We analyze the fundamental sensitivity limits, achievable by the standard (single- and two-arm) optical interferometers using the Gaussian (squeezed coherent) quantum states of the probing light. We consider two types of the measurements of the output light — the standard homodyne measurement and the non-linear threshold measurement [Helstrom, “Quantum detection and estimation theory” (1976)]. The latter one allow to recover in one-shot measurement the phase information carried out by the second momenta of the quadrature amplitudes of the output light, providing the sensitivity saturating the Quantum Cramer-Rao bound.
We show that in all considered cases, the Heisenberg scaling for the phase measurement error $\Delta\phi=K/N$ can be achieved, where $K\sim 1$ is a numerical prefactor and $N$ is the mean photon number. The specific value of the prefactor depends on the interferometer topology, the measurement type, and the quantum state of the probing light. The smaller it is, the more narrow is the range of the values of $\phi$ where the sensitivity that close to the best one for this specific case is provided.
@@ I Introduction
Measurement of the phase of light using optical interferometers is one of the key tasks of experimental physics. Sensitivity of the best modern interferometers is very high and to a major extent is limited by quantum fluctuations of the probing light, see e.g. the review papers [1, 2, 3]. In particular, if the ordinary coherent state of light is used, then the best possible sensitivity corresponds to the Shot Noise Limit (SNL), equal to
@@
where $\Delta\phi$ is the mean square phase measurement error and $N$ is the mean photon number in the probing light. Better sensitivity, for the same value of $N$, can be achieved by using more advanced quantum states. In particular, it was proposed in Ref. [4] to use for this purpose Gaussian quadrature-squeezed states. It was shown in that work that in the case of moderate squeezing, $e^{2r}\ll N$, where $r$ is the logarithmic squeeze factor, the measurement error $\Delta\phi$ can be suppressed by $e^{r}$, giving the following squeezing-enhanced SNL:
@@
The most sensitive contemporary optical interferometers, namely the laser gravitational-wave detectors, like LIGO [5] or VIRGO [6], use this approach. As a result, they can measure relative elongations of their multi-kilometer length arms with the precision of about $\sim\!10^{-23}\,\text{Hz}^{-1/2}$ [7].
In the case of the more exotic non-Gaussian states of light, the sensitivity could reach the Heisenberg limit (HL) [8, 9, 10]:
@@
where $K\sim 1$ is a numerical factor. The same result could be obtained also using Gaussian squeezed states in the case of very strong squeezing, $e^{2r}\sim N$ [11].
@@
Opposite to the well-established limits (1) and (2), the HL is still subject of discussions. Various specific forms of the HL (3) and different values of the parameter $K$ can be found in literature. The main reason for this is the absence of the unique Hermitian phase operator, canonically conjugate to the photon number operator $\hat{N}$ [12]. Different workarounds allowing to overcome this issue were proposed, giving different forms of the HL. In addition, in some works the sensitivity is calculated assuming given mean photon number, while in other ones — assuming given total photon number. Reviews of the various approaches to the HL can be found in Refs. [1, 13, 14].
In Refs. [15, 16, 17], the following form of the HL was obtained using numerical optimization of the phase uncertainty for the given mean photon number $N$:
$\Delta\phi_{\rm HL}=\frac{\approx\!\!1.38}{N+1}\,.$ (4)
In addition, it was shown in Ref. [16], that if the photon number is not very small, $N\gtrsim 5$, then the following analytical solution to this problem can be found:
$\Delta\phi_{\rm HL}=\sqrt{\frac{4}{27}|a_{1}|^{3}}\frac{1}{N+1}\,,$ (5)
were $a_{1}$ being the first zero of the Airy function ${\rm Ai}(\cdot)$, and the quantum state of an oscillator (e.g. the optical mode) that allows to reach this limit is a non-Gaussian one, having, in the Fock representation, the shape close to the Airy function ${\rm Ai}(\cdot)$.
At the same time, it was shown in Ref. [11] that using Gaussian squeezed states and the standard homodyne measurement procedure, it is possible to obtain the sensitivity noticeably exceeding the limit (4):
$\Delta\phi=\frac{1}{2\sqrt{N(N+1)}}$ (6)
This discrepancy could be explained by the fact that the bound (4) is valid for any value of $\phi$. This means that it does not require any a priori information on $\phi$. On the other hand, the limit (6) holds only if $\phi^{2}\lesssim e^{-2r}$. This means that if the sensitivity close to the HL is required, then the a priori uncertainty of the phase must not exceed the SNL value $\sim\!\!1/\sqrt{N}$. The role of the a priori information was explored, in particular, in Refs. [18, 19, 14]. It was shown that it indeed could significantly affect the achievable sensitivity.
@@
In this paper, we analyze the sensitivity limits, achievable by the standard optical interferometers and the Gaussian (squeezed coherent) quantum states. We consider two most important from the practical point of view interferometric topologies, the single-arm and the antisymmetric two-arm ones (see e.g. the review [3]). The conceptually more simple single-arm topology, see Fig. 1, directly implements measurement of the phase shift $\phi$ in a harmonic oscillator mode. In this case, a second beam providing the phase reference is necessary.
@@
In the two-arm configuration, shown in Fig. 2 (the Mach-Zehnder topology is depicted; it is known, however, that the Michelson one is equivalent to it), the phase shifts $\phi$ and $-\phi$ are introduced antisymmetrically into the first and the second arms, respectively, and both beamsplitters are the balanced 50%/50% ones. This configuration is insensitive to the common phase shift and therefore more tolerant to technical noises and drifts. Due to this reason, it is used, in particular, in the GW detectors [20].
We consider two types of the measurements of the output light. The first one, which we use as a reference point, is the well-known homodyne measurement. Its disadvantage is that it is sensitive only to the linear information on $\phi$, that is the one provided by the mean values of the output quadratures of light. The second one is the non-linear threshold measurement. It was introduced in Ref. [21] (see also Sec. IV.3 of the monograph [22]) as the tool for binary discrimination between two given quantum states. Here we adapt it to the continuous measurement task. The advantage of this approach is that it allows to recover the best possible sensitivity, at least for a one given value of $\phi$, while providing the explicit form of the optimal measured variable.
This paper is organized as follows. In Sec. II, we discuss the threshold measurement procedure. In Sections III and IV we calculate the sensitivity of the single-arm and double-arm interferometers, respectively, achievable by using the homodyne detection and the threshold detection. In Sec. V, we propose a possible semi-gedanken implementation of the threshold measurement, based on the quadratic ($\chi^{(2)}$) optical nonlinearity. In Sec. VI, we summarize the obtained results.
II Cramer-Rao bound and the threshold detector
We start with the well-familiar quantum Cramer-Rao bound (QCRB), see Sec. VIII.4 of the monograph [22].
@@ II Cramer-Rao bound and the threshold detector
Let $\hat{\rho}(\phi)$ be the density operator of an object, depending on the parameter $\phi$ that has to be estimated. It can be shown that the lower bound for the variance of an unbiased estimate of $\phi$ exists, which does not depend on the measurement procedure and has the following form:
@@
where ^$\hat{\mathcal{L}}(\phi)$ is the symmetric logarithmic derivative defined by the following equation:
$\dfrac{\partial\hat{\rho}(\phi)}{\partial\phi}=\hat{\rho}(\phi)\circ\hat{\mathcal{L}}(\phi)\,,$ (8)
where here and in the rest of this paper, $\circ$” means symmetrized product, that is, for any operators $\hat{Q}$ and $\hat{P}$,
$\hat{Q}\circ\hat{P}=\frac{1}{2}(\hat{Q}\hat{P}+\hat{P}\hat{Q})\,.$ (9)
It was shown also in Ref. [22] that if the quantum state $\hat{\rho}(\phi)$ is a pure one and has the following form:
@@
^$\hat{\mathcal{R}}(\phi)=e^{-i\hat{\mathcal{N}}\phi}$ (11)
is a unitary displacement operator and $\hat{\mathcal{N}}$ is a Hermitian operator, then Eq. (7) can be presented in the following closed form:
@@
where $(\Delta\mathcal{N})^{2}$ is the variance of $\hat{\mathcal{N}}$ in the state (10). Note that while the operator $\hat{\mathcal{L}}$ explicitly depends on $\phi$, the QCRB (12) does not depend on it.
Now consider the slightly modified (adapted to the measurement of a continous variable) version of the threshold detection concept of Refs. [21, 22]. Suppose that some Hermitian operator $\hat{Y}$ is measured, and we are interested in the sensitivity at some given value of $\phi$, for example, at $\phi=0$ (evidently, this choice does not limit the generality). Using the error propagation approach, the estimation error $\Delta\phi_{0}$, corresponding to the vanishingly small values of $\phi\to 0$ (the threshold sensitivity) can be calculated as follows:
$(\Delta\phi_{0})^{2}=\frac{(\Delta Y_{0})^{2}}{G_{0}^{2}}\,,$ (13)
@@
$(\Delta Y_{0})^{2}=\mathop{\rm Tr}\nolimits[\hat{\rho}(0)\hat{Y}^{2}]-\mathop{\rm Tr}\nolimits[\hat{\rho}(0)\hat{Y}]^{2}$ (14)
is the uncertainty of $\hat{Y}$ at $\phi=0$ and
$G_{0}=\lim_{\phi\to 0}\dfrac{\partial\mathop{\rm Tr}\nolimits[\hat{\rho}(\phi)\hat{Y}]}{\partial\phi}$ (15)
is the gain factor. It can be shown that the minimum of the uncertainty (13) is provided by the operator $\hat{Y}$ satisfying the following equation:
@@
$\mathop{\rm Tr}\nolimits[\rho(0)\hat{Y}]=0$ (17)
and
$G_{0}=(\Delta Y_{0})^{2}\,.$ (18)
Therefore,
$(\Delta\phi_{0})^{2}=\frac{1}{(\Delta Y_{0})^{2}}\,.$ (19)
It is easy to note that up to the notations, for the particular case of $\phi\to 0$ Eq. (8) coincides with Eq. (16). Therefore, the measurement of $\hat{Y}$ saturates the QCRB in this particular case, compare Eqs. (7) and (19). It is natural to expect, that this measurement should provide a good sensitivity, approaching the QCRB, albeit may be not reaching it exactly, for sufficiently small values of $\phi$, satisfying the condition $\phi<\delta\phi$, where $\delta\phi$ is the half-width of this good sensitivity band.
@@ III Single-arm interferometer
Here we consider the single-arm interferometer shown in Fig. 1. First, we calculate the sensitivity, achievable in the case of the homodyne detection. Then we construct the explicit form of the operator $\hat{Y}$ measured by the threshold detector and calculate the sensitivity at $\phi\to 0$. In both cases, we optimize the input state of light, assuming a given mean photon number $N$, and express the phase measurement error in the terms of this $N$. Finally, we calculate the sensitivity, provided by the measurement of $Y$, for the general case of $\phi\neq 0$, using the error propagation method.
We suppose that the probe beam is prepared in the Gaussian squeezed coherent state
@@
$\hat{\mathcal{D}}=e^{\alpha(\hat{a}^{\dagger}-\hat{a})}\,,$ (21)
is the displacement operator (we assume without limiting the generality that the displacement parameter $\alpha$ is real), $\hat{a}$, $\hat{a}^{\dagger}$ are, respectively, the annihilation and creation operators, and
$\hat{\mathcal{S}}=e^{r(\hat{a}^{\dagger}{}^{2}-\hat{a}^{2})/2}\,,$ (22)
is the squeeze operator with the real and non-negative squeeze factor $r$. This choice of $r$ corresponds to the squeezed phase quadrature of the input light and therefore to the best phase sensitivity for a given $r$, see Refs. [4, 3]. It is easy to show that the mean value $N$ and the variance $(\Delta N)^{2}$ of the number of quanta $\hat{N}=\hat{a}^{\dagger}\hat{a}$ in the state (20) are equal to
$N=\alpha^{2}+\mathop{\rm sinh}\nolimits^{2}r\,,$ (23a)
$(\Delta N)^{2}=\alpha^{2}e^{2r}+\frac{1}{2}\mathop{\rm sinh}\nolimits^{2}2r\,.$ (23b)
Homodyne detection.
Using the Heisenberg picture, this scheme can be described by the following input/output relations:
$\begin{pmatrix}\hat{x}_{\rm out}\\ \hat{p}_{\rm out}\end{pmatrix}=\hat{\mathcal{U}}^{\dagger}(\phi)\begin{pmatrix}\hat{x}\\ \hat{p}\end{pmatrix}\hat{\mathcal{U}}(\phi)=\mathbb{R}(\phi)\begin{pmatrix}\sqrt{2}\alpha+\hat{x}e^{r}\\ \hat{p}e^{-r}\end{pmatrix},$ (24)
@@
^$\hat{\mathcal{R}}(\phi)=e^{-i\hat{N}\phi}$ (26)
is the unitary phase shift operator, $\hat{N}$ is the photon number operator,
$\mathbb{R}(\phi)=\begin{pmatrix}\cos\phi&\sin\phi\\ -\sin\phi&\cos\phi\end{pmatrix}$ (27)
is the rotation matrix, $\hat{x}$, $\hat{p}$ are, respectively, the Hermitian amplitude and phase quadrature operators of the input field, defined by
@@
and $\hat{x}_{\rm out}$, $\hat{p}_{\rm out}$ are the corresponding operators of the output field. Note that the squeeze and displacement operations are taken into account in Eqs. (24). Therefore, the operators $\hat{x}$ and $\hat{p}$ corresponds to the initial vacuum field and their variances are equal to $1/2$.
In the case of the small phase shift, $|\phi|\ll 1$, the phase quadrature of the output beam $\hat{p}_{\rm out}$ carries the major part of the phase information. Therefore, consider the homodyne measurement of this quadrature. Using the error propagation method, we obtain that the corresponding measurement error is equal to
@@
In order to present Eq. (29) in the HL-like form, we minimize it in $\alpha^{2}$ under the condition (23a). To find the best possible sensitivity, we perform this optimization at $\phi=0$. In this case, the minimum is achieved at
@@
Threshold detector at $\phi\to 0$
Consider now the threshold detector. In the Schrödinger picture, the output field of the interferometer is described by the following density operator:
where
$\hat{\rho}_{0}=|{\psi_{0}}\rangle\langle{\psi_{0}}|\,.$ (35)
Substituting it into Eq. (16), we obtain:
$i[\hat{\rho}_{0},\hat{N}]=\hat{\rho}_{0}\circ\hat{Y}\,.$ (36)
Consider then a set of quantum states $\{|{\psi_{k}}\rangle\}$ complementing $|{\psi_{0}}\rangle$ to the full orthonormal set:
Using this representation, Eq. (36) can be rewritten as follows:
This equation can be solved explicitly, giving:
$\langle{\psi_{k}}|\hat{Y}|{\psi_{l}}\rangle=\begin{cases}0\,,&k=l=0\,,\\ 2i\langle{\psi_{0}}|\hat{N}|{\psi_{l}}\rangle\,,&k=0\,,\ l>0\,,\\ -2i\langle{\psi_{k}}|\hat{N}|{\psi_{0}}\rangle\,,&k>0\,,\ l=0\,,\\ \text{undefined}\,,&k>0\,,\ l>0\,,\end{cases}$ (39)
It is easy to see from this equation, that
$(\Delta Y_{0})^{2}=4(\Delta N)^{2}$ (40)
and, therefore,
$(\Delta\phi_{0})^{2}=\frac{1}{4(\Delta N)^{2}}\,,$ (41)
where the variance $(\Delta N)^{2}$ is given by Eq. (23b).
Now minimize Eq. (41) under the condition (23a). Taking into account that $\alpha^{2}\geq 0$, it is easy to show that the minimum is provided by
@@
that is, corresponds to the squeezed vacuum state. In this case, it follows from Eqs. (23) that the measurement error at $\phi=0$ is equal to
@@
In order to calculate the measurement error for other values of $\phi$, the explicit form of the operator $\hat{Y}$ is required, that is, the set $\{|{\psi_{k}}\rangle\}$ and the fourth clause in Eq. (39) have to be defined explicitly. In principle, this can be done in infinite number of ways. The sensitivity at $\phi=0$ is fully defined by Eq.(39) and therefore does not depend on this choice, but the sensitivity at other values of $\phi$ could depend. In this paper, we consider a simple bilinear in $\hat{x}$, $\hat{p}$ form of $Y$ that, in principle, could be implemented in an experiment, see Sec. V.
[…]
$\hat{\mathcal{D}}^{\dagger}\hat{a}\hat{\mathcal{D}}=\hat{a}+\alpha\,,$ (45a)
$\hat{\mathcal{S}}^{\dagger}\hat{a}\hat{\mathcal{S}}=\hat{a}\mathop{\rm cosh}\nolimits r+\hat{a}^{\dagger}\mathop{\rm sinh}\nolimits r\,,$ (45b)
we obtain that
and, correspondingly,
$-2i\langle{\psi_{k}}|\hat{N}|{\psi_{0}}\rangle=\langle{k}|\hat{\mathcal{Y}}^{\dagger}|{0}\rangle\,,$ (47)
where
As a result, Eq. (39) takes the following form:
$\langle{k}|\hat{\mathcal{S}}^{\dagger}\hat{\mathcal{D}}^{\dagger}\hat{Y}\hat{\mathcal{D}}\hat{\mathcal{S}}|{l}\rangle=\begin{cases}0\,,&k=l=0\,,\\ \langle{0}|\hat{\mathcal{Y}}|{l}\rangle\,,&k=0\,,\ l>0\,,\\ \langle{k}|\hat{\mathcal{Y}}^{\dagger}|{0}\rangle\,,&k>0\,,\ l=0\,,\\ \text{undefined}\,,&k>0\,,\ l>0\,,\end{cases}$ (49)
A possible form of $\hat{Y}$ that satisfies Eq. (49) for all $k>0,l>0$ is the following:
$\hat{\mathcal{S}}^{\dagger}\hat{\mathcal{D}}^{\dagger}\hat{Y}\hat{\mathcal{D}}\hat{\mathcal{S}}=\hat{\mathcal{Y}}+\hat{\mathcal{Y}}^{\dagger}+\hat{a}^{\dagger}\hat{Q}\hat{a}=-2\hat{p}\circ(\sqrt{2}\alpha e^{r}+\hat{x}\mathop{\rm sinh}\nolimits 2r)+\hat{a}^{\dagger}\hat{\hat{Q}}\hat{a}\,,$ (50)
where $\hat{Q}$ is an arbitrary Hermitian operator. In this paper, we consider the simplest case of $\hat{Q}=0$.
Rolling back the squeeze and displacement operators, we obtain:
$\hat{Y}=\hat{\mathcal{D}}\hat{\mathcal{S}}(\hat{\mathcal{Y}}+\hat{\mathcal{Y}}^{\dagger})\hat{\mathcal{S}}^{\dagger}\hat{\mathcal{D}}^{\dagger}=-2\hat{p}\circ(\sqrt{2}\alpha\mathop{\rm cosh}\nolimits 2r+\hat{x}\mathop{\rm sinh}\nolimits 2r)\,.$ (51)
@@
Now we are in position to calculate the phase measurement error for all values of $\phi$, using the error propagation method. We assume the same value of $\alpha=0$ that gives the best possible sensitivity at $\phi=0$, see Eq. (42). In this case, it follows from Eqs. (23a), (24), and (52) that the mean value and the variance of $\hat{Y}_{\rm out}$ are equal to
@@
$G=\dfrac{\partial\langle{\hat{Y}_{\rm out}}\rangle}{\partial\phi}=8N(N+1)\,.$ (55)
is the gain factor.

### v2

@@ Abstract
We analyze the fundamental limits of phase measurement precision, provided by the standard (single- and two-arm) optical interferometers using the Gaussian (squeezed coherent) quantum states of the probing light. We consider two types of the measurements of the output light — the homodyne measurement and the non-linear threshold measurement that provides the sensitivity saturating the quantum Cramer-Rao bound.
For all considered cases, we calculate the best sensitivity $\Delta\phi_{0}$ achievable at some given value of phase and the range $\delta\phi$ around this value within which the sensitivity is close to $\Delta\phi_{0}$. We show that in all cases, the Heisenberg scaling $\Delta\phi=K/N$ can be reached, where $K\sim 1$ is a numerical prefactor and $N$ is the mean photon number. We show also that $\delta\phi$ strongly depends on $K$.
@@ I Introduction
Measurement of the phase of light using optical interferometers is one of the key tasks of experimental physics. The sensitivity of the best modern interferometers is very high and to a major extent is limited by quantum fluctuations of the probing light, see e.g. the review papers [1, 2, 3]. In particular, if the ordinary coherent state of light is used, then the best possible sensitivity corresponds to the Shot Noise Limit (SNL), equal to
@@
where $\Delta\phi$ is the mean square phase measurement error and $N$ is the mean photon number interacting with the phase shifting object(s). Better sensitivity, for the same value of $N$, can be achieved by using more advanced quantum states. In particular, it was proposed in Ref. [4] to use Gaussian quadrature-squeezed states for this purpose. It was shown in that work that in the case of moderate squeezing, $e^{2r}\ll N$, where $r$ is the logarithmic squeeze factor, the measurement error $\Delta\phi$ can be suppressed by $e^{r}$, giving the squeezing-enhanced SNL:
@@
The most sensitive contemporary optical interferometers, namely the laser gravitational-wave detectors, like LIGO [5] or VIRGO [6], use this approach. As a result, they can measure the relative elongations of their multi-kilometer length arms with precision of about $\sim\!10^{-23}\,\text{Hz}^{-1/2}$ [7].
In the ultimate case of very strong squeezing, $e^{2r}\sim N$, the sensitivity could reach the so-called Heisenberg limit (HL) [8, 2]:
@@
where $K\sim 1$ is a numerical factor. The same result could also be obtained using more exotic non-Gaussian quantum states of light [9, 10, 11, 1, 12]. But, unfortunately, the non-Gaussian states are notoriously sensitive to the optical losses and are very hard to prepare [13].
@@
Opposite to the well-established limits (1) and (2), the HL is still a subject of discussion. Various specific forms of the HL (3) and different values of the parameter $K$ can be found in literature. In addition, in some works the sensitivity is calculated assuming given mean photon number, while in others — assuming given total photon number. Reviews of the various approaches to the HL can be found in Refs. [1, 14, 15].
For example, in Refs. [16, 17, 18, 19], the following sensitivity limit was obtained for the given mean photon number $N$ and an exotic non-Gaussian state:
$\Delta\phi_{\rm HL}\approx\frac{1.38}{N}\,.$ (4)
At the same time, Gaussian squeezed state and the standard homodyne measurement allows to reach the sensitivity noticeably exceeding this value, see Eq (35) below. This discrepancy could be explained by the influence of a priori information on the measurement, see, in particular, Refs. [20, 21, 15]. Note, indeed, that the bound (4) holds for any value of $\phi$. At the same time, the limit (35) can be approached only in a narrow range of the values of $\phi$ with the width $\sim 1/N$ around some given value of the phase, see Eq. (32).
In this paper, we calculate the sensitivity limits, achievable by the standard optical interferometers using the Gaussian (squeezed coherent) quantum states. We take into account a priori information on the phase by calculating the width $\delta\phi$ of the range within which the measurement error does not degrade significantly in comparison to the best possible value $\Delta\phi_{0}$, see Eq. (69). Evidently, if the a priori distribution fits into this range, then the resulting measurement error can be approximated by $\Delta\phi_{0}$.
@@
We consider two most important from the practical point of view interferometric topologies, the single-arm and the antisymmetric two-arm ones (see e.g. the review [3]). The conceptually more simple single-arm topology, see Fig. 1, directly implements measurement of the phase shift $\phi$ in a harmonic oscillator mode.
@@
In the two-arm configuration, shown in Fig. 2 (the Mach-Zehnder topology is depicted; it is known, however, that the Michelson one is equivalent to it), the phase shifts are introduced antisymmetrically into the first and the second arms, respectively, and both beamsplitters are the balanced 50%/50% ones. This configuration is insensitive to the common phase shift and therefore more tolerant to technical noises and drifts. Due to this reason, it is used, in particular, in the GW detectors [22, 23].
We consider two types of specific measurement procedures. The first one is the well-known homodyne detector, which is widely used in high-precision experiments. Yet another commonly used detection procedure is the photon number counting, see e.g. Refs. [24, 1, 25, 26]; we do not consider it here because it gives the results that are close but inferior to the ones of the homodyne detection.
The second procedure which we discuss here is the non-linear threshold detector proposed in Ref. [27] for detection of binary (yes/no) signals. Here we adapt it to the problem of estimation a continuous variable. The key feature of this type of measurement is that for one given value of $\phi$, it automatically saturates the quantum Cramer-Rao bound (QCRB), which gives the fundamental lower bound for the measurement imprecision [28].
This paper is organized as follows. In the introductory Sec. II, we familiarize the readers with the concepts of QCRB and threshold measurement. In Sections III and IV we analyze the single-arm and two-arm interferometers, respectively. In Sec. V, we summarize the obtained results.
II Quantum Cramer-Rao bound and the threshold detector
Quantum Cramer-Rao bound.
@@ II Quantum Cramer-Rao bound and the threshold detector
Let us start with the QCRB. Let $\hat{\rho}(\phi)$ be the density operator of an object, depending on the real parameter $\phi$ (e.g. the phase) that has to be estimated. It can be shown, see Sec. VIII of the monograph Ref. [28], that the lower bound for the variance of an unbiased estimate of $\phi$ exists, which does not depend on the measurement procedure and has the following form:
@@
where $\hat{\mathcal{L}}(\phi)$ is the symmetric logarithmic derivative defined by the following equation:
$\dfrac{\partial\hat{\rho}(\phi)}{\partial\phi}=\hat{\rho}(\phi)\circ\hat{\mathcal{L}}(\phi)\,.$ (6)
Here and in the rest of this paper $\circ$” means symmetrized product, that is, for any operators $\hat{Q}$ and $\hat{P}$,
[…]
We assume in this paper that the quantum state of the probing light is a pure one and has the following form:
@@
$\hat{\mathcal{R}}(\phi)=e^{-i\hat{\mathcal{N}}\phi}$ (9)
it the unitary phase shift operator and $\hat{\mathcal{N}}$ is a Hermitian operator. It was shown in Ref. [28] that in this case Eq. (5) can be presented in the following closed form:
@@
where $(\Delta\mathcal{N})^{2}$ is the variance of $\hat{\mathcal{N}}$ in the state (8).
Threshold detector.
It is known that the measurement of $\mathcal{L}$ saturates the QCRB, see Ref. [28]. Its application to the phase measurement with Gaussian states of light was considered in Ref. [29]. Semi-gedanken implementation of this measurement, based on the second order optical nonlinearity, was proposed in Ref. [30].
Note however that the operator $\hat{\mathcal{L}}$ explicitly depends on $\phi$ (while the QCRB itself does not). This feature creates a “vicious loop”: in order to measure $\phi$, we have to know its value in advance. Therefore, let us consider an adapted to the measurement of a continuous variable version of threshold detection, initially proposed in [27] (see also Sec. IV.3 of [28]). Suppose that some Hermitian operator $\hat{Y}$ is measured, and we are interested in the sensitivity at some given value of $\phi=\phi_{0}$. Without limiting the generality, we assume that $\phi_{0}=0$. Using the error propagation approach, the estimation error $\Delta\phi_{0}$, corresponding to the vanishingly small values of $\phi\to 0$ (the threshold sensitivity) can be calculated as follows:
$(\Delta\phi_{0})^{2}=\frac{(\Delta Y)^{2}}{G^{2}}\biggr|_{\phi\to 0}\,$, (11)
@@
$(\Delta Y)^{2}=\mathop{\rm Tr}\nolimits[\hat{\rho}(\phi)\hat{Y}^{2}]-\bigl(\mathop{\rm Tr}\nolimits[\hat{\rho}(\phi)\hat{Y}]\bigr)^{2}$ (12)
is the variance of $\hat{Y}$ and
$G=\dfrac{\partial\mathop{\rm Tr}\nolimits[\hat{\rho}(\phi)\hat{Y}]}{\partial\phi}$ (13)
is the gain factor.
It can be shown that the minimum of the uncertainty (11) is provided by the operator $\hat{Y}$, which satisfies the following equation:
@@
$\mathop{\rm Tr}\nolimits[\rho(0)\hat{Y}]=0\,,\quad G\bigr|_{\phi=0}=(\Delta Y)^{2}\bigr|_{\phi=0}\,$ (15)
and therefore,
$(\Delta\phi_{0})^{2}=\frac{1}{(\Delta Y)^{2}}\biggr|_{\phi\to 0}\,.$ (16)
It is easy to note that up to the notations, in the case of $\phi\to 0$, Eq. (6) is identical to Eq. (14). Therefore, the measurement of $\hat{Y}$ saturates the QCRB in this particular case, compare Eqs. (5) and (16).
It is natural to expect that for smooth $\hat{\rho}(\phi)$ and for sufficiently small values of $\phi$, this measurement should provide a good sensitivity, approaching the QCRB, albeit may be not reaching it exactly. The corresponding measurement error in this case can be calculated using the generic error propagation formula
$(\Delta\phi)^{2}=\frac{(\Delta Y)^{2}}{G^{2}}\,.$ (17)
This approach will be used in the next sections.
@@ III Single-arm interferometer
Let us consider the single-arm interferometer shown in Fig. 1. We suppose that the input beam is prepared in the Gaussian squeezed coherent state
@@
$\hat{\mathcal{D}}=e^{\alpha(\hat{a}^{\dagger}-\hat{a})}\,$ (19)
is the displacement operator, where we assume without limiting the generality that the displacement parameter $\alpha$ is real, $\hat{a}$, $\hat{a}^{\dagger}$ are, respectively, the annihilation and creation operators, and
$\hat{\mathcal{S}}=e^{r(\hat{a}^{\dagger}{}^{2}-\hat{a}^{2})/2}\,$ (20)
is the squeeze operator with the real and non-negative squeeze factor $r$. This choice of $r$ corresponds to the squeezed phase quadrature of the input light and therefore to the best phase sensitivity for a given $r$, see Refs. [4, 3]. It is easy to show that the mean value $N$ and the variance $(\Delta N)^{2}$ of the number of quanta
$\hat{N}=\hat{a}^{\dagger}\hat{a}$ (21)
in the state (18) are equal to
$N=\alpha^{2}+\mathop{\rm sinh}\nolimits^{2}r\,$, (22)
$(\Delta N)^{2}=\alpha^{2}e^{2r}+\frac{1}{2}\mathop{\rm sinh}\nolimits^{2}2r\,.$ (23)
Using the Heisenberg picture, evolution of optical fields in scheme can be described by the following input/output relations (see e.g. Ref. [3]):
$\begin{pmatrix}\hat{x}_{\rm out}\\ \hat{p}_{\rm out}\end{pmatrix}=\hat{\mathcal{U}}^{\dagger}(\phi)\begin{pmatrix}\hat{x}\\ \hat{p}\end{pmatrix}\hat{\mathcal{U}}(\phi)=\begin{pmatrix}(\sqrt{2}\alpha+\hat{x}e^{r})\cos\phi+\hat{p}e^{-r}\sin\phi\\ -(\sqrt{2}\alpha+\hat{x}e^{r})\sin\phi+\hat{p}e^{-r}\cos\phi\end{pmatrix}$, (24)
@@
$\hat{\mathcal{R}}(\phi)=e^{-i\hat{N}\phi}$ (26)
is the phase shift operator, $\hat{x}$, $\hat{p}$ are, respectively, the Hermitian amplitude and phase quadrature operators of the input field, defined by
@@
and $\hat{x}_{\rm out}$, $\hat{p}_{\rm out}$ are the corresponding operators of the output field. Note that the squeeze and displacement operations are taken into account in Eqs. (24). Therefore, the operators $\hat{x}$ and $\hat{p}$ correspond to the initial vacuum fields and their variances are equal to $1/2$.
Homodyne detection.
In the case of the small phase shift, $|\phi|\ll 1$, the phase quadrature of the output beam $\hat{p}_{\rm out}$ carries the major part of the phase information. Therefore, we consider the homodyne measurement of this quadrature. Using the error propagation method, we obtain that the corresponding measurement error is equal to
@@
If $r\neq 0$, then the variance (29) depends on $\phi$, and if $r>0$, then it increases with the increase of $\phi$, limiting the high sensitivity range $\delta\phi$. Note also that the two terms in parentheses in Eq. (29) are related by the uncertainty principle. Therefore, the stronger the squeezing and therefore the better the sensitivity at $\phi=0$, the bigger the second term. As a result, $\delta\phi$ scales with $r$ as $e^{-2r}$.
In order to present Eq. (28) in HL-like form, we minimize it in $\alpha^{2}$ under the condition (22). To find the best possible sensitivity, we perform this optimization at $\phi=0$. In this case, the minimum is achieved at
@@
In this case, about half of the optical power is provided by the coherent amplitude and the other part — by the squeezing, see Eq. (22). It is interesting that the same result was obtained in Ref. [24] for the photon number counting case.
@@
QCRB and threshold detector at $\phi\to 0$
It follows from Eqs. (10), (9), and (26) that the QCRB in the single-arm interferometer case is equal to
$(\Delta\phi)^{2}=\frac{1}{4(\Delta N)^{2}}\,$, (33)
where the photon number variance is given by Eq. (23). Minimizing it in $\alpha^{2}$ under the condition (22) and taking into account that $\alpha^{2}\geq 0$, we obtain that the minimum is provided by
@@
and is equal to
@@
In order to calculate the measurement error for other values of $\phi$, the explicit form of the operator $\hat{Y}$ is required. It is shown in App. A that it can be presented in the following bilinear form, compare with Refs. [31, 32, 29]:
$\hat{Y}=-2\hat{p}\circ(\sqrt{2}\alpha\mathop{\rm cosh}\nolimits 2r+\hat{x}\mathop{\rm sinh}\nolimits 2r)\,.$ (36)
@@
Now we are in a position to calculate the phase measurement error for all values of $\phi$, using the error propagation method. We assume the same value of $\alpha=0$ that gives the best possible sensitivity at $\phi=0$, see Eq. (34). In this case, it follows from Eqs. (22), (24), and (37) that the mean value and the variance of $\hat{Y}_{\rm out}$ are equal to
@@
$G=\dfrac{\partial\langle{\hat{Y}_{\rm out}}\rangle}{\partial\phi}=8N(N+1)\cos 2\phi\,.$ (40)